# Joint Pricing and Resource Allocation: An Optimal Online-Learning Approach

### v1

@@ 1 Introduction
On the other hand, the problem of resource allocation – to serve different demand classes with various types of resources – presents a complex challenge within the field of operations research. Analogous to online dynamic pricing, the recent proliferation of e-platforms has magnified the importance of developing online allocation algorithms that efficiently manage supply and demand on the fly while maximizing cumulative utilities. However, traditional approaches in resource allocation are insufficient in depicting scenarios where the demand is stochastic and dependent on the price. Therefore, it is critical to develop price-dependent online allocation models and methodologies that can simultaneously learn the demand curve and optimize the joint decisions on price, inventory, and allocation.
@@
As described above, we assume that the inventories are perishable and the leftover inventory cannot be carried over to the following period. Furthermore, we assume that the price is identical for all consumers. We formalize the above process as solving the following two-stage stochastic programming problem:
@@
Here $g(\vec{I},p,\vec{D})$ is the minimum negative net profit under the best allocation scheme given inventories $\vec{I}$, price $p$, and realized demand $\vec{D}$ (see Eq. 2 for its definition).
@@
In this work, we establish a novel framework of solving online pricing and allocation problem under demand uncertainties. Our main contributions are threefold:
@@
    Algorithmic Design against Non-Convexity. We propose an efficient online-learning algorithm for the (price,inventory) joint decision problem, which is highly non-convex. To navigate to the global optimal decisions among many sub-optimals, our algorithm incorporates an optimistic meta-algorithm with multiple online convex optimization (OCO) agents working locally.
@@
    Closed-Form Allocation Scheme. We present an algorithm that outputs a price-dependent closed-form solution to the optimal allocation problem.
  3. 3.
    Regret Analysis. Our algorithm achieves $\tilde{O}(\sqrt{Tmn})$ regret, which is optimal with respect to $T$ as it matches the information-theoretic lower bound in (Broder and Rusmevichientong,, 2012).
1.2 Technical Novelty.
@@
Confidence bound in vertical and horizontal directions.
According to our problem setting, we suffer from bandit feedback and non-convexity with respect to price $p$. Existing methodologies in dynamic pricing mostly adopt continuum-armed bandits to resolve the challenges, nonetheless leading to $O(T^{2/3})$ regret that is sub-optimal in our setting. In this work, we design an $O(\sqrt{T})$-regret algorithm, which makes a careful and novel use of the piecewise-convex property of the objective function. Specifically, we assign an agent to work on each convex interval, conducting online convex optimization (OCO) in local domains. Each agent maintains two high-probability confidence bounds: (1) A vertical bound on the value of local optimal, updated at the end of each sub-epoch. (2) A horizontal bound on the location of local optimal, updated at the end of each epoch.
@@
In order to distinguish among every local-optimality iterated by a local OCO agent, we develop an LCB meta-algorithm over those agents. With the help of the vertical bound of each agent, we estimate the least possible value in each convex interval within reasonable error, and select the most optimistic agent (i.e., the one with the least lower bound) to make decisions in the following sub-epoch.
Closed-form solution of the second-stage LP.
In the second stage when inventories are allocated to fulfill the realized demands, we solve a linear program (LP) that maximizes the net profit under the corresponding demand and supply constraints. In this work, we develop an efficient polynomial-time algorithm that achieves a closed-form solution to this LP. (On the contrary, an LP is generally not guaranteed to have closed-form solutions.) We also show the dependence of this solution on the price $p$, from which we derive the piecewise-convex property of the objective function.
@@ 3 Preliminaries
In this section, we rigorously define the problem we are studying. We firstly formulate the offline version of the problem as a two-stage stochastic program in Section 3.1. The properties of its solution are presented in Section 3.2. We then develop the formulation of the online version in Section 3.3, where the demand parameters are unknown. Finally, we present assumptions that are crucial to our online algorithm design by the end of this section.
@@ 3.1 Offline Problem Setting
We consider the following scenario where a retail company makes their decisions with the goal of maximizing its net profit. Suppose the company produces and sells identical products, and it has $m$ warehouses and $n$ retailers. In general, this company faces the following three problems on inventory, pricing and allocation:
@@
    How to allocate inventories from warehouses to stores in the face of heterogeneous supply costs?
@@
    In Stage 1, the company makes inventory decisions $\vec{I}=[I_{1},I_{2},\ldots,I_{m}]^{\top}$, where $I_{i}$ represents the inventory level of Warehouse $i$. Each unit of inventory at Warehouse $i$ incurs an inventory cost $\gamma_{i}$.
    In addition, the company decides a uniform price $p$ for the products.
@@
The company aims to make the best (inventory, price) decisions that maximize their net profit, which can be formulated as the following optimization problem (where we equivalently minimize the negative net profit):
@@
$\quad\quad\sum_{i=1}^{m}X_{i,j}\leq D_{j},\forall j\in[n]$
$\quad\quad\sum_{j=1}^{n}X_{i,j}\leq I_{i},\forall i\in[m].$
@@
We denote an optimal solution to Eq. 1 as $(p^{*},\vec{I}^{*})$, and denote an optimal solution to Eq. 2 as $\mathbf{X}^{*}$. Since Eq. 2 is a linear program, we may directly solve it with any off-the-shelf optimization tool. However, in order to solve Eq. 1, we have to know the distribution of $\vec{D}$ and how it is dependent on the price $p$, both of which are not directly accessible from the seller’s side as they do not have the full knowledge of the entire market. In the next subsection, we will discuss how we can “learn” the demand distributional function under mild assumptions.
3.2 Solution to $g(\vec{I},p,\vec{D})$
Before we study the online problem, we first provide a solution to the optimization problem in Eq. 2 and investigate the properties of $g(\vec{I},p,\vec{D})$. We firstly propose a greedy algorithm to solve Eq. 2.
Algorithm 1 Greedy $g(\vec{I},p,\vec{D})$ solver
1:  Input: Parameters $C_{i,j},i\in[m],j\in[n]$; Variable $\vec{I}\in\mathbb{R}^{m},p\in\mathbb{R}^{+},\vec{D}\in\mathbb{R}^{n}$.
2:  Sort $\{(i,j)\}_{i=1,j=1}^{i=m,j=n}$ tuples into $\{(i_{k},j_{k})\}_{k=1}^{mn}$ according to Algorithm 2.
3:  Denote $C_{i_{0},j_{0}}:=0$ and $C_{i_{mn+1},j_{mn+1}}:=p_{\max}$ as a complementary definition of $\{C_{i_{K},j_{K}}\}$.
4:  Find $K\in\{0\}\cup[mn]$ such that
$C_{i_{K},j_{K}}<p\leq C_{i_{K+1},j_{K+1}}$ (3)
5:  Let $\tilde{I}\leftarrow\vec{I},\tilde{D}\leftarrow\vec{D},\mathbf{X}\leftarrow 0^{m\times n}$.
6:  for Epoch $k=1,2,3,\ldots,K$ do
7:     Set $X_{i_{k},j_{k}}\leftarrow\min\{\tilde{I}_{i_{k}},\tilde{D}_{j_{k}}\}$.
8:     Update $\tilde{I}_{i_{k}}\leftarrow\tilde{I}_{i_{k}}-X_{i_{k},j_{k}}$.
9:     Update $\tilde{D}_{j_{k}}\leftarrow\tilde{D}_{j_{k}}-X_{i_{k},j_{k}}$.
10:  end for
11:  RETURN $\mathbf{X}$.
@@ 3.2 Solution to $g(\vec{I},p,\vec{D})$
We show the correctness of Algorithm 1 as follows.
Theorem 3.1.
The matrix $\mathbf{X}$ returned by Algorithm 1 is an optimal solution to the optimization problem in Eq. 2.
While the complete proof of Theorem 3.1 is relegated to Section A.1, we provide a high-level idea of the proof below. Notice that Eq. 2 presents a linear program (LP), and therefore strong duality holds. The proof is conducted in the following steps:
  1. 1.
    Construct the dual problem of Eq. 2.
  2. 2.
    Propose an algorithm that solves this dual problem and propose a feasible dual solution.
  3. 3.
    Show that the value of this dual solution equals that of the primal solution provided by Algorithm 1.
Next, we show the marginal convexity of $g(\vec{I},p,\vec{D})$.
Lemma 3.2.
The key to proving Lemma 3.2 is to show that $g(\vec{I},p,\vec{D})$ is the piecewise maximum over a group of linear functions. Please refer to Section A.2 for details.
3.3 Online Problem Setting
Due to insufficient knowledge of the actual demand distribution, the company might propose $(p,\vec{I})$ pairs that are suboptimal, leading to lower net profits compared to the optimal solution. However, the company has observations on the realized demand at each store, which enables them to estimate demand and subsequently enhance their decision-making. In what follows, we study the online decision-making problem of setting prices and managing inventory.
Assumption 3.3.
Assume the realized demand is linear and noisy. Specifically, assume
$\vec{D}_{t}=\vec{a}-\vec{b}\cdot p+\vec{N}_{t}.$ (4)
Based on the definition of functions $Q$ and $Q_{t}$ above, we know that $Q_{t}(\vec{I},p)$ (and therefore $Q(\vec{I},p)$) is marginally convex on $\vec{I}$, since $g(\vec{I},p,\vec{D})$ is marginally convex on $\vec{I}$ (according to Lemma 3.2) and $\langle\vec{\gamma},\vec{I}\rangle$ is linear. Also, we have the following properties.
Lemma 3.4.
Denote $C_{i_{0},j_{0}}=0$ and $C_{i_{mn+1},j_{mn+1}}=p_{\max}$. For any $K\in\{0\}\cup[mn]$, function $Q_{t}(\vec{I},p)$ is Lipschitz and marginally convex on $p$ in range $[C_{i_{K},j_{K}},C_{i_{K+1},j_{K+1}}]$.
We defer the proof of Lemma 3.4 to Section A.3. Furthermore, we get following results (in a non-trivial way).
@@ 3.3 Online Problem Setting
$W(p)$ is Lipschitz. We denote its Lipschitz coefficient as $L_{W}$ for future use. Also, for any $K\in\{0\}\cup[mn]$, function $W(p)$ is convex in range $[C_{i_{K},j_{K}},C_{i_{K+1},j_{K+1}}]$.
The proof of Lemma 3.5 is relegated to Section A.4. Finally, we define the regret as the relative loss on net profit compared to that achieved by optimal decisions.
@@
At each time $t=1,2,\ldots,T$, denote $\vec{I}_{t}$ and $p_{t}$ as the inventory and price decisions, respectively. Define
@@
Before we conclude this section, we present some assumptions that are crucial to our algorithm design.
@@
We have full knowledge on the problem parameters $\vec{\gamma},\{C_{i,j}\}_{i=1,j=1}^{m,n}$ and the boundedness parameters $\gamma_{\max},a_{\max},b_{\max},I_{\max},p_{\max}$ before $t=0$. We have no prior knowledge on the model parameters $\vec{a},\vec{b}$ nor any information about the distribution of $\vec{N}_{t}$.
@@ 4 Algorithm
1:  Input: $m,n,T$, sorted supply costs $\{C_{i_{k},j_{k}}\}_{k=0}^{mn+1}$.
2:  Initialization: Agents $\mathcal{A}_{K}$ (defined in Algorithm 4) and parameters $\hat{W}_{K}:=0,\Delta_{K}:=+\infty,LCB_{K}:=-\infty,K=0,1,2,\ldots,mn$, global parameters $\delta_{K},n_{0}$.
3:  while $t\leq T$ do
4:     Let $\hat{K}:=\mathop{\mathrm{argmin}}_{K\in[mn]\cup 0}LCB_{K}$.
5:     Run agent $\mathcal{A}_{\hat{K}}$ for one sub-epoch, and get updated $\hat{W}_{K}$ and $\Delta_{K}$
6:     Update $LCB_{K^{\prime}}\leftarrow\hat{W}_{K^{\prime}}-34\cdot\Delta_{K^{\prime}}$ for all $K^{\prime}=0,1,2,\ldots,mn.$
7:  end while
Algorithm 4 Optimization Agent $K$
1:  Input: Price interval $[C_{i_{K},j_{K}},C_{i_{K+1},j_{K+1}}]$, parameter $\delta_{K},n_{0},L_{W}$.
2:  Output: $\hat{W}_{K},\Delta_{K}$ (at the end of each sub-epoch)
3:  Initialization: $L_{K,1}=C_{i_{K},j_{K}},U_{K,1}=C_{i_{K+1},j_{K+1}},\hat{W}_{K}=0,\Delta_{K}=+\infty$, and $\qquad a_{K,1}=\frac{3L_{K,1}+U_{K,1}}{4},c_{K,1}=\frac{L_{K,1}+U_{K,1}}{2},b_{K,1}=\frac{L_{K,1}+3U_{K,1}}{4}.$
4:  for $p=a_{K,1},b_{k,1},c_{k,1}$do
5:     Propose inventory $\vec{I}_{0}:=[1,1,\ldots,1]^{\top}$ and price $p$.
6:     Record the marginal function of $\vec{I}$ as $Q_{0}(\vec{I},p)$.
7:     Find $I_{K,1,0}(p):=\mathop{\mathrm{argmin}}_{\vec{I}}Q_{0}(\vec{I},p)$.
8:  end for
9:  STAGE 1: Search for optimality asymptotically
10:  (See Algorithm 5 for details)
11:  STAGE 2: Complementary sampling to enhance $\Delta_{K}$
12:  (See Algorithm 6 for details)
13:  STAGE 3: Pure local-exploitation
14:  (See Algorithm 7 for details)
Algorithm 5 Agent $\mathcal{A}_{K}$ Stage 1
1:  for Epoch $\tau=1,2,\ldots,O(\log T)$do
2:     Let $a_{K,\tau}=\frac{3L_{K,1}+U_{K,1}}{4},c_{K,\tau}=\frac{L_{K,1}+U_{K,1}}{2},b_{K,\tau}=\frac{L_{K,1}+3U_{K,1}}{4}$.
3:     for  Sub-epoch $s=1,2,\ldots$ do
4:        Let sub-epoch length $n_{s}:=2^{s}$
5:        Define a $flag:=0$ for error-bar update.
6:        for $\hat{p}_{\tau}=a_{K,\tau},b_{K,\tau},c_{K,\tau}$ do
7:           for $t=1,2,\ldots,n_{s}$ do
8:              Propose decisions $(\vec{I}_{t},p_{t})=(\vec{I}_{K,\tau,s-1}(\hat{p}_{\tau}),\hat{p}_{\tau})$.
9:              Observe and record the marginal function $Q_{t}(\vec{I},\hat{p}_{\tau})$ with respect to $\vec{I}$.
@@
11:           Define an aggregated function $Q_{K,\tau,s}(\vec{I},\hat{p}_{\tau}):=\frac{1}{n_{s}}\cdot\sum_{t=1}^{n_{s}}Q_{t}(\vec{I},\hat{p}_{\tau})$.
12:           Define the empirical optimal inventory $\qquad I_{K,\tau,s}(\hat{p}_{\tau}):=\mathop{\mathrm{argmin}}_{\vec{I}}Q_{K,\tau,s}(I,\hat{p}_{\tau})$.
13:           Denote $\hat{Q}_{K,\tau,s,\hat{p}_{\tau}}:=Q_{k,\tau,s}(I_{K,\tau,s}(\hat{p}_{\tau}),\hat{p}_{\tau})$, and $\Delta_{K,\tau,s}:=\frac{\delta_{K}}{2\sqrt{n_{s}}}$.
14:        end for
15:        if $\hat{Q}_{K,\tau,s,a_{K,\tau}}>\hat{Q}_{K,\tau,s,b_{K,\tau}}+4\Delta_{K,\tau,s}$ then
16:           Update $L_{K,\tau+1}\leftarrow a_{K,\tau}$, $U_{K,\tau+1}\leftarrow U_{K,\tau}$, $flag\leftarrow 1$.
17:        else if $\hat{Q}_{K,\tau,s,a_{K,\tau}}<\hat{Q}_{K,\tau,s,b_{K,\tau}}-4\Delta_{K,\tau,s}$ then
18:           Update $L_{K,\tau+1}\leftarrow L_{K,\tau}$, $U_{K,\tau+1}\leftarrow b_{K,\tau}$, $flag\leftarrow 1$.
19:        else if $\hat{Q}_{K,\tau,s,c_{K,\tau}}<\hat{Q}_{K,\tau,s,a_{K,\tau}}-4\Delta_{K,\tau,s}$ then
20:           Update $L_{K,\tau+1}\leftarrow a_{K,\tau}$, $U_{K,\tau+1}\leftarrow U_{K,\tau}$, $flag\leftarrow 1$.
21:        else if $\hat{Q}_{K,\tau,s,c_{K,\tau}}<\hat{Q}_{K,\tau,s,b_{K,\tau}}-4\Delta_{K,\tau,s}$ then
22:           Update $L_{K,\tau+1}\leftarrow L_{K,\tau}$, $b_{K,\tau+1}\leftarrow U_{K,\tau}$, $flag\leftarrow 1$.
23:        end if
24:        if $flag==1$ then
25:           if $U_{K,\tau+1}-L_{K,\tau+1}>\frac{1}{T}$ then
26:              Continue to Epoch $\tau+1$ (without updating $\hat{W}_{K}$ or $\Delta_{K}$).
27:           else
28:              Set $\hat{p}_{K}^{*}\leftarrow C_{K,\tau},\hat{\vec{I}}_{K,0}^{*}\leftarrow\vec{I}_{K,\tau,s}(C_{K,\tau})$.
29:              Break to STAGE 2 (without updating $\hat{W}_{K}$ or $\Delta_{K}$).
30:           end if
31:        else if $\Delta_{K,\tau,s-1}<\Delta_{K}$ then
32:           Update $\Delta_{K}\leftarrow\Delta_{K,\tau,s-1},\hat{W}_{K}\leftarrow\min_{\hat{p}_{\tau}\in\{a_{K,\tau},c_{K,\tau},b_{K,\tau}\}}\hat{Q}_{K,\tau,s-1,\hat{p}_{\tau}}$.
33:        end if
34:     end for
35:  end for

### v2

@@ 1 Introduction
On the other hand, the problem of resource allocation – to serve different demand classes with various types of resources – presents a complex challenge within the field of operations research. Analogous to online dynamic pricing, the recent proliferation of e-platforms has magnified the importance of developing online allocation algorithms that efficiently manage supply and demand on the fly while maximizing cumulative utilities. However, traditional approaches in resource allocation are insufficient in depicting scenarios where the demand is stochastic and dependent on the price (Hwang et al.,, 2021). Therefore, it is critical to develop price-dependent online allocation models and methodologies that can simultaneously learn the demand curve and optimize the joint decisions on price, inventory, and allocation.
@@
We assume that the inventories are perishable and the leftover inventory cannot be carried over to the following period. Furthermore, we assume that the price is identical for all consumers. We formalize the above process as solving the following two-stage stochastic programming problem:
@@
Here $g(\vec{I},p,\vec{D})$ is the minimum negative net profit (or loss) under the best allocation scheme given inventories $\vec{I}$, price $p$, and realized demand $\vec{D}$ (see Eq. 2 for its definition).
@@
In this work, we establish a novel framework for solving the online pricing and allocation problem under demand uncertainties. Our main contributions are twofold:
@@
    Algorithmic Design against Non-Convexity. We propose an efficient online-learning algorithm for the (price,inventory) joint decision problem, which is highly non-convex. To navigate to the global optimal decisions among many sub-optimals, our algorithm incorporates an optimistic meta-algorithm to manage multiple online convex optimization (OCO) agents working locally.
@@
    Regret Analysis. We show that our algorithm achieves $\tilde{O}(\sqrt{Tmn})$ regret, which is optimal with respect to $T$ as it matches the information-theoretic lower bound in (Broder and Rusmevichientong,, 2012).
1.2 Statement of Novelty.
@@
Confidence Bounds: Local (Vertical) and Interval (Horizontal) Perspectives
Our setting poses two main challenges: bandit feedback and non-convexity with respect to price $p$. Existing dynamic pricing methods address these with continuum-armed bandit algorithms, but these typically yield $O(T^{2/3})$ regret (Kleinberg,, 2004), which is not optimal in this scenario. To improve upon this, we leverage the piecewise-convex structure of the expected loss: The domain $[0,p_{\max}]$ can be divided into intervals, in each of which the function is locally convex. We assign a dedicated agent to each interval, applying online convex optimization (OCO) to approach its local optimal. We introduce twofold confidence bounds for each agent:
  1. (a)
    Vertical confidence bound: This quantifies the uncertainty in the agent’s estimate of the local optimal value within its convex interval. It is updated at the end of each sub-epoch. (We provide the definitions of sub-epoch and epoch in Section 4).
  2. (b)
    Horizontal confidence bound: This describes the shrinking search space within the interval where the local optimum is likely to be found. It is updated at the end of each epoch.
As the algorithm proceeds, either the vertical confidence bound (value uncertainty) or the horizontal confidence bound (location uncertainty) for each agent becomes tighter, thereby improving both the accuracy of local optimization and the efficiency of interval selection. This hierarchical design enables us to achieve an improved regret bound of $O(\sqrt{T})$.
@@
In order to distinguish among every local optimum, we develop an LCB meta-algorithm over these local OCO agents. With the vertical confidence bound of each agent, we obtain the least possible value in each interval and select the agent with the least lower bound to run in the following sub-epoch.
@@ 3 Preliminaries
In this section, we rigorously define the problem we are studying. We firstly formulate the offline version of the problem as a two-stage stochastic program in Section 3.1. We then develop the formulation of the online version in Section 3.2, where the demand parameters are unknown. Finally, we present assumptions that are crucial to our online algorithm design by the end of this section.
@@ 3.1 Offline Problem Setting
We consider the following scenario where a retail company makes their decisions with the goal of maximizing their net profit. Suppose that the company has $m$ warehouses and $n$ retailers, producing and selling identical products. In general, this company faces the following three problems on inventory, pricing, and allocation:
@@
    How to allocate inventories from warehouses to stores under heterogeneous supply costs?
@@
    In Stage 1, the company makes inventory decisions $\vec{I}=[I_{1},I_{2},\ldots,I_{m}]^{\top}$, where $I_{i}$ is the inventory level of Warehouse $i$. Each unit of inventory at Warehouse $i$ incurs an inventory cost $\gamma_{i}$. In addition, the company decides a uniform price $p$ for the products.
@@
The company aims to make the best (inventory, price) joint decisions that maximize their net profit, which can be formulated as the following optimization problem (where we equivalently minimize the negative net profit):
@@
$\quad\quad\sum_{i=1}^{m}X_{i,j}\leq D_{j},\forall j\in[n];\quad\sum_{j=1}^{n}X_{i,j}\leq I_{i},\forall i\in[m].$
@@
We denote an optimal solution to Eq. 1 as $(p^{*},\vec{I}^{*})$, and denote an optimal solution to Eq. 2 as $\mathbf{X}^{*}$. Since Eq. 2 is a linear programming, we can solve it directly with any standard optimization tool. However, in order to solve Eq. 1, we have to know the distribution of $\vec{D}$ and how it is dependent on the price $p$, both of which are not directly accessible from the seller’s side as they do not have the full knowledge of the entire market. In the next subsection, we will discuss how we can “learn” the demand distribution function under mild assumptions.
Now we propose a lemma that states the marginal convexities of $g(\vec{I},p,\vec{D})$.
Lemma 3.1.
The key to proving Lemma 3.1 is to show that $g(\vec{I},p,\vec{D})$ is the piecewise maximum over a group of linear functions. Please refer to Section B.1 for details.
3.2 Online Problem Setting
Due to insufficient knowledge of the actual demand distribution, the company could propose pairs of $(p,\vec{I})$ that are suboptimal, leading to lower net profits compared to the optimal solution. However, the company has observations on the realized demand at each store, which enables them to estimate demand and subsequently improve their decision-making. In what follows, we study the online decision-making problem of setting prices and managing inventory.
Assumption 3.2.
Assume the realized demand is linear and noisy. Specifically, assume
$\vec{D}_{t}=\vec{a}-\vec{b}\cdot p+\vec{N}_{t}.$ (3)
Since $g(\vec{I},p,\vec{D})$ is marginally convex on $\vec{I}$ according to Lemma 3.1, we know that $Q(\vec{I},p)$ and $Q_{t}(\vec{I},p)$ are also marginally convex on $\vec{I}$. But what about their marginal behaviors on $p$? We state in the following lemma:
Lemma 3.3.
Sort $\{C_{i,j}\}_{i=1,j=1}^{m,n}$ according to Algorithm 1. Denote $C_{i_{0},j_{0}}=0$ and $C_{i_{mn+1},j_{mn+1}}=p_{\max}$. For any $K\in\{0\}\cup[mn]$, function $Q_{t}(\vec{I},p)$ is Lipschitz and marginally convex on $p$ in range $[C_{i_{K},j_{K}},C_{i_{K+1},j_{K+1}}]$.
@@ 3.2 Online Problem Setting
We defer the proof of Lemma 3.3 to Section B.2. Furthermore, we have the following results.
@@
We have $W(p)$ is $L_{W}$-Lipschitz where $L_{W}$ is a constant. Also, for any $K\in\{0\}\cup[mn]$, the function $W(p)$ is convex in the range $[C_{i_{K},j_{K}},C_{i_{K+1},j_{K+1}}]$.
The proof of Lemma 3.4 is relegated to Section B.3. Finally, we define regret as the relative loss of net profit compared to that achieved by optimal decisions.
@@
At each time $t=1,2,\ldots,T$, denote $\vec{I}_{t}$ and $p_{t}$ as inventory and price decisions, respectively. Define
@@
Before we conclude this section, we present some crucial assumptions for our algorithm design.
@@
We have full knowledge on the problem parameters $\vec{\gamma},\{C_{i,j}\}_{i=1,j=1}^{m,n}$ and the boundedness parameters $\gamma_{\max},a_{\max},b_{\max},I_{\max},p_{\max}$ before $t=0$. We do not know the model parameters $\vec{a},\vec{b}$ nor the distribution of $\vec{N}_{t}$.
@@ 4 Algorithm
1:  Input: $m,n,T$, sorted supply costs $\{C_{i_{k},j_{k}}\}_{k=0}^{mn+1}$. Global parameters $\delta_{K},n_{0}$.
2:  Initialization: $\hat{W}_{K}:=0,\Delta_{K}:=+\infty,LCB_{K}:=-\infty,K=0,1,2,\ldots,mn$.
3:  for $K=0,1,2,\ldots,mn$ do
4:     Initialize local agent $\mathcal{A}_{K}$ as follows:
5:     Let $L_{K,1}=C_{i_{K},j_{K}},U_{K,1}=C_{i_{K+1},j_{K+1}},$and $a_{K,1}=\frac{3L_{K,1}+U_{K,1}}{4},c_{K,1}=\frac{L_{K,1}+U_{K,1}}{2},b_{K,1}=\frac{L_{K,1}+3U_{K,1}}{4}.$
6:     for $p=a_{K,1},b_{k,1},c_{k,1}$do
7:        Propose inventory $\vec{I}_{0}:=[1,1,\ldots,1]^{\top}$ and price $p$.
8:        Record the marginal function of $\vec{I}$ as $Q_{0}(\vec{I},p)$.
9:        Find $I_{K,1,0}(p):=\mathop{\mathrm{argmin}}_{\vec{I}}Q_{0}(\vec{I},p)$.
@@
11:     Set the Stage Flag £$\pounds_{K}\leftarrow 1$ for Agent $\mathcal{A}_{K}$.
12:  end for
13:  while $t\leq T$ do
14:     Let $\hat{K}:=\mathop{\mathrm{argmin}}_{K\in[mn]\cup 0}LCB_{K}$.
15:     if £$\pounds_{\hat{K}}==1$ then
16:        Run Algorithm 3 (Stage 1 of $\mathcal{A}_{\hat{K}}$) for one sub-epoch.
[…]
18:        Update $\hat{W}_{\hat{K}},\Delta_{\hat{K}}$ and £$\pounds_{\hat{K}}$ according to the statement of $\mathcal{A}_{\hat{K}}$.
19:     else if £$\pounds_{\hat{K}}==2$ then
20:        Run Algorithm 4 (Stage 2 of $\mathcal{A}_{\hat{K}}$) until its completion.
21:        $t+=|\text{the length of this Stage 2}|$.
22:        Update $\hat{W}_{\hat{K}},\Delta_{\hat{K}}$ according to the statement of $\mathcal{A}_{\hat{K}}$.
23:        Update £$\pounds_{\hat{K}}\leftarrow 3$.
24:     else if £$\pounds_{\hat{K}}==3$ then
25:        Run Algorithm 5 (Stage 3 of $\mathcal{A}_{\hat{K}}$) for one single time period.
26:        $t+=1$.
27:        Update $\hat{W}_{\hat{K}},\Delta_{\hat{K}}$ according to the statement of $\mathcal{A}_{\hat{K}}$.
28:     end if
29:     for $K^{\prime}=0,1,2,\ldots,mn.$ do
30:        Update $LCB_{K^{\prime}}\leftarrow\hat{W}_{K^{\prime}}-34\cdot\Delta_{K^{\prime}}$.
31:     end for
32:  end while